OpenStudy (anonymous):

find a quadratic polynomial each with the given numbers as the sum and products of its zeroes respectively.....1/4,-1

Directrix (directrix):

A quadratic in the form of ax2+bx +c = 0 has sum of roots: −b/a and product of roots: c/a. Sum of roots: ========= -b/a = 1/4 -a = 4b a = -4b ------- Product of roots: c/a = -1/1 c = -a --------- Let b = -1. Then a = 4 and because c = -a, then c = -4. The coefficients of a, b, and c are: 4, -1, and -4 or a multiple of those 3. 4x^2 -1 x - 4 = 0

OpenStudy (anonymous):

2 months ago? and it's still open....

OpenStudy (anonymous):

For fuzzy bunnies and giggles, here's a better answer: For a polynomial \(P(x)\), we have \(P(x)=(x-r_1)(x-r_2)=x^2-(r_1+r_2)x+r_1r_2\). Therefore, our polynomial is \(x^2-\frac{1}{4}x-1=4x^2-x-4\).